Simultaneous Equations
3x+2y=16
x+2y=12

Answer by Alan3354(69443)

Simultaneous Equations
3x+2y=16
x+2y=12
===================
That's what they are.
If you want to solve them, you should say that.
3x+2y=16
x +2y=12
----------------- Subtract
2x = 4
x = 2
y = 5